# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v2

@@ Abstract
The Internet of Medical Things (IoMT) transcends traditional medical boundaries, enabling a transition from reactive treatment to proactive prevention. This innovative method revolutionizes healthcare by facilitating early disease detection and tailored care, particularly in chronic disease management, where IoMT automates treatments based on real-time health data collection. Nonetheless, its benefits are countered by significant security challenges that endanger the lives of its users due to the sensitivity and value of the processed data, thereby attracting malicious interests. Moreover, the utilization of wireless communication for data transmission exposes medical data to interception and tampering by cybercriminals. Additionally, anomalies may arise due to human error, network interference, or hardware malfunctions. In this context, anomaly detection based on Machine Learning (ML) is an interesting solution, but it comes up against obstacles in terms of explicability and privacy protection. To address these challenges, a new framework for Intrusion Detection Systems (IDS) is introduced, leveraging Artificial Neural Networks (ANN) for intrusion detection while utilizing Federated Learning (FL) for privacy preservation. Additionally, eXplainable Artificial Intelligence (XAI) methods are incorporated to enhance model explanation and interpretation. The efficacy of the proposed framework is evaluated and compared with centralized approaches using multiple datasets containing network and medical data, simulating various attack types impacting the confidentiality, integrity, and availability of medical and physiological data. The results obtained offer compelling evidence that the FL method performs comparably to the centralized method, demonstrating high performance. Additionally, it affords the dual advantage of safeguarding privacy and providing model explanation while adhering to ethical principles.
@@ 1 INTRODUCTION
To mitigate these concerns regarding anomalies, ML has been proposed as a solution. Considering that IoMT systems generate large amounts of data, they can help ML models to distinguish between normal and abnormal behaviour. This capability facilitates the detection of anomalies and is effective against zero-day and new attacks. These detection systems can operate in real-time, and with advancements in Deep Learning (DL), the process of attribute selection and image processing [3] becomes automated.
@@
From an ethical perspective, the proposed framework addresses critical concerns such as fairness, accountability, and bias mitigation. By leveraging FL, the framework ensures that sensitive data remains on local devices, reducing the risk of biased outcomes that can arise from centralized data collection. This decentralized approach allows for the inclusion of diverse datasets, promoting fairness and reducing the likelihood of discriminatory results. Furthermore, the integration of XAI enhances the transparency and interpretability of the ML models, enabling stakeholders—including patients, model designers, and regulators—to understand how decisions are made. This transparency fosters accountability, as it becomes easier to identify and address potential biases or errors in the system. By providing clear and understandable explanations of the model’s predictions, XAI ensures compliance with regulatory requirements that mandate explainability, such as those outlined in HIPAA, GDPR, WHO and ISO.
Through transparent explanations, a proven track record of reliable performance, and the provision of detection history in percentage form, this framework cultivates trust among users, facilitating confidence in the predictive capabilities of the system. By offering intuitive and accessible insights, even non-technical users can access results and track the model’s performance over time. The ethical design of the framework ensures that it not only meets technical and regulatory standards but also aligns with societal values, promoting the responsible and equitable use of AI in healthcare and beyond.
@@
The article is structured as follows: Section 2 presents the background and reviews related work. Section 3 details the methodology for the proposed IDS model. Section 4 describes the experimental setup and presents the results. Section 5 compares the proposed framework with prior studies. Section 6 discusses the contributions and implications of the work. Finally, Section 7 concludes the study and outlines future research directions.
@@ 2.1 Background
In the context of automating the surveillance and analysis of events within an information system, the IDS plays a pivotal role. Categorized based on the source of information they rely on, Network-based IDS (NIDS) focuses on monitoring network traffic, while Host-based IDS (HIDS) centers its analysis on activities occurring on individual hosts. The methods of analysis employed by IDS vary, with Misuse Detection generating alerts when an event matches a predefined signature, and anomaly detection utilizing ML to learn normal behavior and triggering alerts upon detecting deviations [5].
Shifting to ML, which can be categorized into supervised, unsupervised, and semi-supervised learning paradigms. Supervised ML involves training models on labeled data for classification or regression tasks, capturing relationships between inputs and output data [6]. Unsupervised learning explores patterns in unlabeled data [7], while semi-supervised learning leverages both labeled and unlabeled datasets to improve model performance and generalization [8]. The response time of an IDS spans the spectrum from real-time detection to a predefined time interval. Upon detecting an attack, the system initiates a response, which may be passive, involving the notification of the administrator regarding the presence of an attack, or active, entailing actions like communication blockage. The architecture can take different forms, such as Centralized, Fully-Distributed, and Partially Distributed IDS with hierarchical reporting mechanisms.
Introducing FL, a distributed ML approach that preserves privacy by transmitting models instead of raw data. In FL, a global model is constructed by leveraging the participation of multiple devices over communication rounds. A subset of clients is chosen at the beginning of each round, updating the model independently using locally stored data. The server aggregates these models to create an enhanced global model, iterating until convergence. Two types of FL emerge: Cross-Silo FL, involving distributed data centers, and Cross-Device FL, encompassing mobile devices and IoT equipment [9]. FL encompasses various categorizations depending on the characteristics of the data distribution. These include Horizontal FL, where data across multiple devices share identical features but possess distinct samples; Vertical FL, where data across diverse entities exhibit differing features but may share common identifiers; and Federated Transfer Learning, which extends conventional Transfer Learning to the federated setting, allowing the transfer of knowledge across distributed data sources. [10], each contributing to data privacy and learning process improvement.
@@
$TP$ True Positives
$FP$ False Positives
$TN$ True Negatives
$FN$ False Negatives
$IoMT$ Internet of Medical Things
$FL$ Federated Learning
$XAI$ eXplainable Artificial Intelligence
$IDS$ Intrusion Detection System
$HIPAA$ Health Insurance Portability and Accountability Act
$WHO$ World Health Organization
$GDPR$ General Data Protection Regulation
$ISO$ International Organization for Standardization
$IEC$ International Electrotechnical Commission
$ML$ Machine Learning
$DL$ Deep Learning
$ANN$ Artificial Neural Network
$DNN$ Deep Neural Network
$SVM$ Support Vector Machine
$RF$ Random Forest
$KNN$ K-Nearest Neighbors
$CNN$ Convolutional Neural Network
$LSTM$ Long Short-Term Memory
$XSS$ Cross-Site Scripting
$TTL$ Time To Live
@@
$ACCS$ Australian Centre for Cyber Security
$EHMS$ Enhanced Healthcare Monitoring System
$EMR$ Electronic Medical Record
$EHR$ Electronic Health Record
$CISO$ Chief Information Security Officer
$SHAP$ Shapley Additive Explanations
$MITM$ Man-In-The-Middle
$CSV$ Comma Separated Values
$AUC$ Area Under the Curve
$ROC$ Receiver Operating Characteristic
$LIME$ Local Interpretable Model-Agnostic Explanations
$GRU$ Gated Recurrent Unit
$PCA$ Principal Component Analysis
$AE$ Auto Encoder
$LRGU$ Logistic Redundancy Coefficient Upweighting
$MIFS$ Mutual Information Feature Selection
$GWO$ Grey Wolf Optimization
$SGD$ Stochastic Gradient Descent
$MCPS$ Medical Cyber-Physical Systems
$DoS$ Denial of Service
$DDoS$ Distributed Denial of Service
$HFL$ Hierarchical FL
$SMOTE$ Synthetic Minority Oversampling Technique
$DNS$ Domain Name System
$SGRU$ Sliced Gated Recurrent Unit
$sSAE$ Stacked Sparse Autoencoder
$ICU$ Intensive Care Unit
$TCP$ Transmission Control Protocol
$UDP$ User Datagram Protocol
$FTP$ File Transfer Protocol
$SSH$ Secure Shell
$HTTP$ Hypertext Transfer Protocol
$MLP$ Multi-Layer Perceptron
$LR$ Logistic Regression
$RFE$ Recursive Feature Elimination
$NB$ Naive Bayes
$PSO$ Particle Swarm Optimization
$NIDS$ Network-based IDS
$HIDS$ Host-based IDS
$DT$ Decision Tree
$SRU$ Simple Recurrent Unit
$Dintpkt$ Destination Inter Packet
$dstjitter$ Destination Jitter
$dstload$ Destination Load
$srcload$ Source Load
$dst\_port$ destination ports
$src\_port$ source port
However, ML models are often considered black-box. To address this, Van Lent introduced XAI in 2004 [11]. XAI comprises methodologies aiding researchers in comprehending and gaining trust in ML model outcomes. A comprehensive AI model can be characterized with a three-dimensional description [12]. Explainability refers to the capacity to articulate the learning model’s processes, interpretability provides insight into the model’s operation, and transparency denotes inherent understandability without user intervention. In the domain of explainability, transparent models, like Decision Tree (DT), embody ante-hoc explainability, while opaque models, like Deep Neural Network (DNN), require post-hoc explainability. Nevertheless, there are scenarios where even transparent models with seemingly straightforward rules require post-hoc elucidation. Despite the complexity of opaque models, they often outperform transparent ones, leading to a trade-off between performance and explainability. Post-hoc methods operate on both local and global scop, employing techniques such as feature relevance analysis, which assesses the significance of each input feature in the model’s predictions. Surrogate models contribute to interpretability by simplifying the primary model through methods like probing with local changes or utilizing its structure. Additionally, representative examples, drawn from the model’s training set, showcase significant levels of confidence in their classification towards a particular class [13].

### v3

@@ Abstract
The Internet of Medical Things (IoMT) transcends traditional medical boundaries, enabling a transition from reactive treatment to proactive prevention. This innovative method revolutionizes healthcare by facilitating early disease detection and tailored care, particularly in chronic disease management, where IoMT automates treatments based on real-time health data collection. Nonetheless, its benefits are countered by significant security challenges that endanger the lives of its users due to the sensitivity and value of the processed data, thereby attracting malicious interests. Moreover, the utilization of wireless communication for data transmission exposes medical data to interception and tampering by cybercriminals. Additionally, anomalies may arise due to human error, network interference, or hardware malfunctions. In this context, anomaly detection based on Machine Learning (ML) is an interesting solution, but it comes up against obstacles in terms of explicability and privacy protection. To address these challenges, a new framework for Intrusion Detection Systems (IDS) is introduced, leveraging Artificial Neural Networks (ANN) for intrusion detection while utilizing Federated Learning (FL) for privacy preservation. Additionally, eXplainable Artificial Intelligence (XAI) methods are incorporated to enhance model explanation and interpretation. The efficacy of the proposed framework is evaluated and compared with centralized approaches using multiple datasets containing network and medical data, simulating various attack types impacting the confidentiality, integrity, and availability of medical and physiological data. The results offer compelling evidence that the FL method performs comparably to the centralized method, demonstrating high performance. Additionally, it affords the dual advantage of safeguarding privacy and providing model explanation while adhering to ethical principles.
@@ 1 INTRODUCTION
To mitigate these concerns regarding anomalies, ML is proposed as a solution. Considering that IoMT systems generate large amounts of data, they can help ML models to distinguish between normal and abnormal behaviour. This capability facilitates the detection of anomalies and is effective against zero-day and new attacks. These detection systems can operate in real-time, and with advancements in Deep Learning (DL), the process of attribute selection and image processing [3] becomes automated.
@@
Furthermore, the integration of XAI enhances the transparency and interpretability of the ML models, enabling stakeholders—including patients, model designers, and regulators—to understand how decisions are made. By providing clear and understandable explanations of the model’s predictions, XAI ensures compliance with regulatory requirements that mandate explainability, such as those outlined in HIPAA, GDPR, WHO and ISO. Through transparent explanations, a proven track record of reliable performance, and the provision of detection history in percentage form, this framework cultivates trust among users.
From an ethical perspective, the proposed framework addresses critical concerns such as fairness, accountability, and bias mitigation. By leveraging FL, the framework ensures that sensitive data remains on local devices, reducing the risk of biased outcomes that can arise from centralized data collection. This decentralized approach allows for the inclusion of diverse datasets, promoting fairness and reducing the likelihood of discriminatory results. The transparency provided by XAI fosters accountability, as it becomes easier to identify and address potential biases or errors in the system. By offering intuitive and accessible insights, even non-technical users can access results and track the model’s performance over time. The ethical design of the framework ensures that it not only meets technical and regulatory standards but also aligns with societal values, promoting the responsible and equitable use of AI in healthcare and beyond.
@@
The article is structured to provide a logical progression from foundational concepts to empirical validation and contextual analysis. Section 2 first establishes the essential background on IDS, ML, FL, and XAI, followed by a review of related work to identify specific research gaps in ML-based security for IoMT systems. Building on this foundation, Section 3 details the methodology for the proposed IDS model, beginning with a description of the IoMT system architecture and the local training process. It then elaborates on the integrated use of a FL process for secure, privacy-compliant aggregation and XAI techniques for interpretability, an architecture designed to inherently address the ethical imperatives of data protection, transparency, and accountability. Section 4 validates the framework through a multi-faceted evaluation, first examining the impact of FL parameters, then applying SHAP for model interpretability, and finally comparing performance against a centralized baseline. To position the contribution within the field, Section 5 compares the framework’s effectiveness and advancements with prior studies. Broader implications, limitations, and insights derived from the results are synthesized in Section 6. Section 7 concludes by summarizing key findings and outlining promising directions for future research.
@@ 2.1 Background
The automation of surveillance and analysis within information systems critically relies on IDS. These systems are categorized by their data source: Network-based IDS (NIDS) monitor network traffic, while Host-based IDS (HIDS) analyze activities on individual hosts. Their analytical methods are broadly divided into two paradigms: Misuse Detection, which generates alerts by matching events against a database of predefined attack signatures, and Anomaly Detection, which leverages ML to model normal system behavior and flag significant deviations from it [5]. Architecturally, IDS can be implemented as Centralized, Fully-Distributed, or Partially-Distributed systems with hierarchical reporting mechanisms.
ML provides the foundational algorithms that empower modern anomaly-based detection. Standard ML paradigms include supervised learning, which trains models on labeled datasets to perform classification or regression tasks; unsupervised learning, which explores inherent patterns and structures within unlabeled data [6]; and semi-supervised learning, which leverages both labeled and unlabeled data to improve model performance and generalization where labeled data is scarce [7]. The application of these techniques is particularly salient for the IoMT, which generates vast volumes of complex and heterogeneous data. The capacity of ML to learn directly from data and adapt to evolving environments makes it highly suitable for detecting novel and sophisticated cyber threats in healthcare networks, a task where traditional static rule-based systems often prove inadequate.
FL has emerged as a distributed ML paradigm that directly addresses the core IoMT challenge of data privacy. In FL, a global model is iteratively trained by aggregating model updates from multiple participants without ever transferring or centralizing raw, sensitive patient data [8]. FL architectures are categorized based on the participant type into Cross-Silo FL, involving a few organizational entities, and Cross-Device FL, involving a large number of IoT devices. Furthermore, FL systems are classified based on data partitioning into Horizontal FL, where data across devices share the same feature space but concern different samples; Vertical FL, where data across entities contain different features but may overlap in sample identifiers; and Federated Transfer Learning, which extends transfer learning techniques to facilitate knowledge transfer across distributed data sources with limited common features or samples [9]. For IoMT, FL offers a significant advantage by enabling collaborative training on decentralized data silos, thus complying with stringent data protection regulations and facilitating the development of robust, generalized intrusion detection models that learn from a broader and more diverse data landscape than any single institution possesses.
@@
TPTP True Positives
FPFP False Positives
TNTN True Negatives
FNFN False Negatives
IoMTIoMT Internet of Medical Things
FLFL Federated Learning
XAIXAI eXplainable Artificial Intelligence
IDSIDS Intrusion Detection System
HIPAAHIPAA Health Insurance Portability and Accountability Act
WHOWHO World Health Organization
GDPRGDPR General Data Protection Regulation
ISOISO International Organization for Standardization
IECIEC International Electrotechnical Commission
MLML Machine Learning
DLDL Deep Learning
ANNANN Artificial Neural Network
DNNDNN Deep Neural Network
SVMSVM Support Vector Machine
RFRF Random Forest
KNNKNN K-Nearest Neighbors
CNNCNN Convolutional Neural Network
LSTMLSTM Long Short-Term Memory
XSSXSS Cross-Site Scripting
TTLTTL Time To Live
@@
ACCSACCS Australian Centre for Cyber Security
EHMSEHMS Enhanced Healthcare Monitoring System
EMREMR Electronic Medical Record
EHREHR Electronic Health Record
CISOCISO Chief Information Security Officer
SHAPSHAP Shapley Additive Explanations
MITMMITM Man-In-The-Middle
CSVCSV Comma Separated Values
AUCAUC Area Under the Curve
ROCROC Receiver Operating Characteristic
LIMELIME Local Interpretable Model-Agnostic Explanations
GRUGRU Gated Recurrent Unit
PCAPCA Principal Component Analysis
AEAE Auto Encoder
LRGULRGU Logistic Redundancy Coefficient Upweighting
MIFSMIFS Mutual Information Feature Selection
GWOGWO Grey Wolf Optimization
SGDSGD Stochastic Gradient Descent
MCPSMCPS Medical Cyber-Physical Systems
DoSDoS Denial of Service
DDoSDDoS Distributed Denial of Service
HFLHFL Hierarchical FL
SMOTESMOTE Synthetic Minority Oversampling Technique
DNSDNS Domain Name System
SGRUSGRU Sliced Gated Recurrent Unit
sSAEsSAE Stacked Sparse Autoencoder
ICUICU Intensive Care Unit
TCPTCP Transmission Control Protocol
UDPUDP User Datagram Protocol
FTPFTP File Transfer Protocol
SSHSSH Secure Shell
HTTPHTTP Hypertext Transfer Protocol
MLPMLP Multi-Layer Perceptron
LRLR Logistic Regression
RFERFE Recursive Feature Elimination
NBNB Naive Bayes
PSOPSO Particle Swarm Optimization
NIDSNIDS Network-based IDS
HIDSHIDS Host-based IDS
DTDT Decision Tree
SRUSRU Simple Recurrent Unit
DintpktDintpkt Destination Inter Packet
dstjitterdstjitter Destination Jitter
dstloaddstload Destination Load
srcloadsrcload Source Load
dst_portdst\_port destination ports
src_portsrc\_port source port
Despite their high predictive efficacy, complex ML and FL models often operate as black boxes, creating a significant barrier to their adoption in high-stakes clinical settings. XAI addresses this critical issue by providing methodologies to make model decisions understandable, interpretable, and trustworthy [10]. A comprehensive framework for XAI can be described across three dimensions: explainability as the capacity to articulate the model’s internal processes, interpretability as the ability to provide intuitive insight into the model’s reasoning, and transparency as the inherent understandability of the model’s mechanics [11]. Models range from being intrinsically interpretable, such as Decision Trees (DT), to opaque, requiring post-hoc explanation techniques, such as Deep Neural Networks (DNN). This often creates a trade-off, where the superior performance of opaque models must be balanced against the need for explainability. Post-hoc methods operate on both local and global levels, employing techniques like feature relevance analysis, surrogate models, and representative examples to elucidate the model’s behavior [12]. In the context of IoMT security, XAI is not a luxury but a necessity. It allows security analysts and healthcare professionals to audit model predictions, verify that alerts are based on clinically relevant features rather than data artifacts, and build the trust required for operational deployment. This capability is crucial for ensuring that automated decisions do not jeopardize patient safety and that the reasoning behind any alert can be rigorously validated, which is paramount for regulatory compliance and clinical adoption.